Please help!! I just started algebra 11/2 week ago so I am new at this..
Find the value of n when (7.17 + n) + 8.83= 16+4

Answer by checkley71(8403)

7.17+N+8.83=16+4 OR N=20-7.17-8.83 OR N=4